Calculate the yield of CO2 on combustion of propane gas (based on - propane+O2 gives CO2 and water), starting with 10g of propane and 4g of O2. Show your calculations
Answer #1

Reaction is

C3H8 + 5 O2 ...............> 3 CO2 + 4 H2O

one mole propane react with 5 mole O2 to form 3 mole CO2 and 4 mole H2O.

10 g of propane = mass / molar mass = 10 g / 44 g / mole = 0.2273 mole.

and

4 g of O2 = 4 g / 32 g / mole = 0.125 mole.

propane is the limiting reactant.

thus

yield of CO2 on combustion of propane gas = 3 * 0.2273 = 0.6819 mole

or

= mole * molar mass

or

= 0.6819 mole * 44 g / mole = 30.0 g

thus

yield of CO2 on combustion of propane gas = 30.0 g (answer)